OpenStudy (anonymous):

Find the equation of the line containing the given pair of points. The equation of the line in slope-intercept form is y= These are the point (-7,-2) (-3,-1)

OpenStudy (anonymous):

x-4y=1 is the equation of the line.

OpenStudy (anonymous):

So that is what y=?

OpenStudy (anonymous):

y=1/4 x -1/4

OpenStudy (anonymous):

y=mx=b are you using that formula? ook

OpenStudy (anonymous):

y=mx+b is what I meant!

OpenStudy (radar):

I think he used m=(-2-(-1))/(-7-(-3)) m=-1/-4=1/4

OpenStudy (anonymous):

m=1/4 and the b = -1/4

OpenStudy (anonymous):

so, the equation is y = 1/4 x - 1/4

OpenStudy (radar):

Collegegirl did you unders stand how redepsilon got the slope from those two points?

OpenStudy (radar):

\[m=\Delta y/\Delta x\]